OpenStudy (anonymous):

Identify the conic section that matches the equation shown below. Then, identify the center and foci. 5x2 + 7y2 + 50x - 28y + 118 = 0

OpenStudy (anonymous):

it is an ellipse, you can tell because you have \(x^2\) and \(y^2\), with both coefficients positive, but not equal

OpenStudy (anonymous):

ok, that makes sense...

OpenStudy (anonymous):

you want to write it in the form \[\frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\] so you can see the center it will be \((h,k)\)

OpenStudy (anonymous):

this is a pain and requires completing the square do you know how to do it?

OpenStudy (anonymous):

wait let me look at my notes a sec...

OpenStudy (anonymous):

how would i do it if theres an x and an x^2

OpenStudy (anonymous):

that is why you have to complete the square

OpenStudy (anonymous):

so.. separate the x's and the y's or something right?

OpenStudy (anonymous):

\[5x^2 + 7y^2 + 50x - 28y + 118 = 0\] \[5x^2+50x+7x^2-28x=-118\] \[5(x^2+10x)+7(x^2-4x)=-118\] \[5(x+5)^2+7(x-2)^2=-118+5\times 5^2+7\times 2^2\] and so on

OpenStudy (anonymous):

but you changed the y after 28 to x...

OpenStudy (anonymous):

actually you can see the center from there it will be \((-5,2)\) but you still need to do more work to get it in the form that you want, so you can find the focus yes, that was a typo on my part

OpenStudy (anonymous):

\[5(x+5)^2+7(y-2)^2=-118+5\times 5^2+7\times 2^2\]

OpenStudy (anonymous):

ohh i see, so all that in the parenthesis after 7 is y

OpenStudy (anonymous):

\[5(x+5)^2+7(y-2)^2=35\] divide both sides by \(35\) and you will get exactly the form that you want

OpenStudy (anonymous):

so the center would be (-5,2)?

OpenStudy (anonymous):

and the foci?

OpenStudy (anonymous):

what did you get for the equation?

OpenStudy (anonymous):

in any case you should have \[\frac{(x+5)^2}{7}+\frac{(y-2)^2}{5}=1\] and so \(a^2=7,b^2=5\) and \[\sqrt{a^2-b^2}=\sqrt{7-5}=\sqrt{2}\]

OpenStudy (anonymous):

so the foci are \(\sqrt{2}\) units to the left and right of the center \((-5,2)\)

OpenStudy (anonymous):

ok my equation was off but i see why noww, and i got the came center

OpenStudy (anonymous):

look right for a foci point?

OpenStudy (anonymous):

\[2\pm \sqrt{2}, -5\]